hartnn (hartnn):

2 easy questions 1) what does this notation mean ?? \(\Large (\dfrac{\partial u}{\partial x})_y \) 2) will ask later in comments.

hartnn (hartnn):

i have u=lx+my,v=mx-ly and i need to prove \(\Large (\dfrac{\partial u}{\partial x})_y \times \Large (\dfrac{\partial x}{\partial u})_v = \dfrac{l^2 }{l^2+m^2}\) proving part will be easier once i know what that notation means!

hartnn (hartnn):

also i already know \(\Large (\dfrac{\partial u}{\partial x})\) means differentiating 'u' partially w.r.t x

hartnn (hartnn):

would it be \(\Large \dfrac{\partial }{\partial y} \Large (\dfrac{\partial u}{\partial x})\) ?

hartnn (hartnn):

And my 2nd Question was : Solve : \(x^6-x^5+x^4-x^3+x^2-x+1 =0 \) I think i need to use DeMovire's Theorem, just need a start. (or any other method that does not involve use of calculator)

ganeshie8 (ganeshie8):

\[\large \left(\dfrac{\partial x}{\partial u}\right)_v \] paritual u keeping v constant ( y depends on v)

ganeshie8 (ganeshie8):

\(u=lx+my, \\ v=mx-ly\) from 1st equation : \[\large u = lx + m\dfrac{(mx-v)}{l} \] \[\large u = x(l + m^2/l) - mv/l \] \[\large 1 = (l + m^2/l) \left(\dfrac{\partial x}{\partial u}\right)_v - 0 \] \[\large \dfrac{l}{l^2+m^2} = \left(\dfrac{\partial x}{\partial u}\right)_v \]

ganeshie8 (ganeshie8):

u,v,x,y are not independent, they depend on each other.. the subsript(v) makes it explicit what variable we're keeping constant

hartnn (hartnn):

nice! got it :) how about 2nd question ?

ganeshie8 (ganeshie8):

@SithsAndGiggles

ganeshie8 (ganeshie8):

the second question looks hard...

hartnn (hartnn):

a similar example in notes was \(x^7+64x^4+64x^3+64^2=0\) but that was easily factorable...this isn't...

OpenStudy (anonymous):

Notice that \[x^6-x^5+x^4-x^3+x^2-x+1=\sum_{n=0}^6(-x)^n\] There's a formula for finite geometric series, but it escapes me at the moment...

OpenStudy (anonymous):

I think it's something like \[\sum_{n=0}^6(-x)^n=\frac{1-(-x)^{n+1}}{1-(-x)}\]

OpenStudy (anonymous):

So when setting this equal to zero, it's the same as finding the 7th roots of \(x^7=-1\), excluding \(x=-1\) itself because that would make the expression undefined.

ganeshie8 (ganeshie8):

brilliant xD

hartnn (hartnn):

so i just prove ( x^6-x^5+x^4-x^3+x^2-x+1)(1+x) = x^7+1

hartnn (hartnn):

which would be easy and then exclude x=-1

OpenStudy (anonymous):

Yeah pretty much :)

hartnn (hartnn):

nice! any other method comes to mind ?

OpenStudy (anonymous):

Newton's method? or any approximation technique.

OpenStudy (anonymous):

Though I doubt that work due to the complex answers.

hartnn (hartnn):

well then, would use your method...thanks! :)

OpenStudy (anonymous):

yw!